Nomber of h+ ions preaent in 250ml of lemon juice of pH=3 is

Grade:12

1 Answers

Arun
6 years ago
From basic pH formula
 
3 = - log[H+]
Now
[H+] = 10^-3
Now 1000 ml = 10^-3 moles
 
Hence
In 250 ml = (1/4) * 10^-3 moles
Hence
Number of H + ions = 6.023 * 10^23 * (1/4) * 10^-3
 = 1.50575 * 10^20 ions
 
Regards
Arun (askIITians forum expert)